Sawyer is 6 yrs older than his sister. This year, Sawyer is as three times older than his sister. How old is Sawyer now?
Answer by macston(5194)

S=Sawyers's age; S-6=sister's age
S=3(S-6)
S=3S-18 Add 18 to each side
S+18=3S Subtract S from each side
18=2S divide each side by 2
9=S ANSWER: Sawyer is 9 years old
CHECK
Sister is (S-6)=3 years old
Sawyer is 3 times sister's age
9 years=3(3 years)
9 years=9 years